### v3

# Out of the Memory Barrier: A Highly Memory Efficient Training System for LLMs with Million-Token Contexts

Abstract
Training Large Language Models (LLMs) on long contexts is severely constrained by prohibitive GPU memory overhead, not training time. The primary culprits are the activations, whose memory footprints scale linearly with sequence length. We introduce OOMB, a highly memory-efficient training system that directly confronts this barrier. Our approach employs a chunk-recurrent training framework with on-the-fly activation recomputation, which maintains a constant activation memory footprint ($\mathcal{O}(1)$) and shifts the primary bottleneck to the growing KV cache. To manage the KV cache, OOMB integrates a suite of synergistic optimizations: a paged memory manager for both the KV cache and its gradients to eliminate fragmentation, asynchronous CPU offloading to hide data transfer latency, and page-level sparse attention to reduce both computational complexity and communication overhead. The synergy of these techniques yields exceptional efficiency. Our empirical results show that for every additional 10K tokens of context, the end-to-end training memory overhead increases by a mere 10MB for Qwen2.5-7B. This allows training Qwen2.5-7B with a 4M-token context on a single H200 GPU, a feat that would otherwise require a large cluster using context parallelism. This work represents a substantial advance in resource efficiency for long-context LLM training. The source code is available at https://github.com/wenhaoli-xmu/OOMB.
1 Introduction
Figure 1: An overview of the OOMB training framework. (Left) OOMB processes sequences chunk-by-chunk with activation recomputation, maintaining a constant activation memory footprint ($\mathcal{O}(1)$) and shifting the bottleneck to the KV cache. (Center) The growing KV cache is managed through a combination of paged memory, asynchronous CPU offloading, and page-level sparse attention. (Right) Performance benchmarks show that the integrated system drastically reduces peak GPU memory, enabling efficient training on long contexts with minimal memory overhead.
Long-context modeling remains a formidable challenge in training Large Language Models (LLMs). The principal obstacle is not only training time but also the prohibitive GPU memory overhead associated with long sequences (Liu et al., 2025b; Wang et al., 2024). Research indicates that significant long-context capabilities can be achieved with a surprisingly small number of training steps (Peng et al., 2024; Ding et al., 2024), and that fine-tuning on as little as 5% of the original pretraining data may suffice for commercial-grade performance (Gao et al., 2025; Bai et al., 2024).
[…]
While various training-free methods for context extension exist (Jin et al., 2024; Han et al., 2024), their practical utility is limited. Prior work has established that common evaluation metrics like perplexity and targeted retrieval are unreliable indicators of true long-range reasoning (Gao et al., 2024; Bai et al., 2024). Achieving robust performance necessitates dedicated contiguous pretraining.
To this end, this paper directly confronts the central problem: how to train longer-context models with fewer resources without compromising performance. We introduce Out Of the Memory Barrier (OOMB), a chunk-recurrent training framework that processes sequences in segments. During the forward pass, each chunk’s activations are computed and then immediately discarded. For the backward pass, they are recomputed on-the-fly. This strategy maintains a constant activation memory footprint, regardless of the total sequence length. This approach, however, shifts the primary bottleneck to the KV cache, which must be retained throughout the training step and scales linearly with context length.
To overcome this challenge, OOMB integrates a system of co-designed optimizations for efficient KV cache management:
  • Paged KV Cache and Gradient Management. We introduce a paged memory manager, inspired by (Kwon et al., 2023), for both the KV cache and its gradients. This design eliminates expensive memory operations and mitigates fragmentation when appending new key-value pairs.
  • Specialized Kernels. We implement custom operators that execute all operations on the KV cache and its gradients directly, making them opaque to PyTorch’s autograd system. This avoids storing the KV cache as an activation and allows for direct gradient accumulation.
  • Asynchronous KV Cache CPU Offloading. We designed an asynchronous mechanism to preemptively offload the KV cache and its gradients, whose memory grows with context length, to CPU memory. The resulting data transfer latency is effectively masked by computation.
  • Page-level Sparse Attention. Our paged memory architecture natively supports page-level sparse attention (Yuan et al., 2025; Lu et al., 2025). This reduces the computational complexity of attention and minimizes the communication overhead of our offloading mechanism.
The synergy of these techniques yields exceptional GPU memory efficiency. Our results show that for every additional 10K context tokens, the end-to-end training memory overhead increases by a mere 10MB for Qwen2.5-7B. This allows training Qwen2.5-7B with a 4-million-token context on a single H200 GPU. In contrast, context parallelism methods (Liu et al., 2024; Lin, 2025; Shoeybi et al., 2020) require large clusters for similar tasks (Xu et al., 2025), highlighting a substantial advance in resource efficiency.
Table 1: Comparison of Long-Context Training Methods. OOMB achieves $\mathcal{O}(1)$ activation memory complexity, where NN denotes the context length. This key advantage allows OOMB, when combined with other techniques, to practically enable single-GPU training on contexts exceeding 4M tokens.
Method GPUs Activation Exact Attn Max Tokens
SeCO (Li et al., 2025) $\mathcal{O}(1)$ 128K
[…]
ZeRO3 Offload $\mathcal{O}(N)$ 128K
Ring Flash Attention $\mathcal{O}(N/8)$ 256K
[…]
OOMB + Dense Attn $\mathcal{O}(1)$ 4M
[…]
2 Related Work
Efficient Long-Context Extension. A core challenge in extending language models to long contexts is their failure to generalize to positional encodings not seen during pretraining (Han et al., 2024). While continued training can resolve this issue, the associated GPU memory costs are often prohibitive. This has motivated training-free context extension methods, which typically modify positional encodings to accommodate longer sequences (Chen et al., 2023; Han et al., 2024; Jin et al., 2024; An et al., 2024). However, these approaches introduce significant trade-offs. Strategies based on interpolation or extrapolation of positional encodings can degrade performance on short-text tasks by reducing the resolution of positional information (Gao et al., 2024). Others that reuse position indices achieve high scores on perplexity benchmarks but have been shown to fail on tasks requiring deep contextual understanding, indicating a superficial grasp of the extended context (Gao et al., 2024).
Given these limitations, state-of-the-art models still rely on fine-tuning to expand their context windows (Grattafiori et al., 2024; Yang et al., 2025; Liu et al., 2025a). While recent advances in parameter-efficient fine-tuning (PEFT) have successfully minimized the memory footprint of model weights and optimizer states (Zhao et al., 2025; Zhang et al., 2025; Zhong et al., 2025), they do not alleviate the activation memory bottleneck that scales linearly with sequence length. Consequently, the prevailing trend has shifted from algorithmic efficiency (Chen et al., 2024) toward overcoming memory barriers with massive computational resources (Liu et al., 2024; Li et al., 2024; Jacobs et al., 2024; Shoeybi et al., 2020). For instance, recent efforts have required a 256-GPU cluster to extend a model’s context window to four million tokens (Xu et al., 2025).
Serial vs. Parallel Training Paradigms. Parallel training processes an entire sequence in a single pass, maximizing GPU utilization and scalability. However, its memory footprint scales linearly with sequence length, creating a significant bottleneck for long-context models. In contrast, serial training is highly memory-efficient, as it only activates a small part of the network at any given time, but at the cost of reduced parallelism.
For long-context fine-tuning, where dataset sizes are often moderate, memory consumption rather than throughput is the primary constraint (Peng et al., 2024; Gao et al., 2024). This observation motivates a hybrid approach that processes sequences in chunks: parallel within each chunk and serial between them. This strategy is standard in modern LLM inference engines (Kwon et al., 2023; Ye et al., 2025) where it incurs negligible latency. Its application to training, however, remains underdeveloped. An early exploration, SeCO (Li et al., 2025), introduced chunk-wise training but lacked critical optimizations at the operator and memory management levels. Consequently, it could only train a 16K context model on a single consumer-grade GPU. Under identical conditions, our fully-optimized system extends this capability by more than tenfold.
Context Parallelism. Context Parallelism (CP) and our proposed method both segment long sequences but diverge significantly in their scalability and overhead profiles. CP, inspired by Ring Attention (Liu et al., 2024) and implemented in frameworks like Megatron-LM (Shoeybi et al., 2020) and DeepSpeed Ulysses (Jacobs et al., 2024), distributes a single sequence across multiple GPUs. This design introduces substantial inter-GPU communication overhead that scales with the number of devices. In contrast, the overhead of our serial approach is limited to repeated KV cache accesses and the latency from CPU offloading. The trade-off is clear: CP excels with large GPU clusters and specialized high-bandwidth networks, whereas our method maintains a moderate and predictable overhead, eliminating the need for such extensive infrastructure.
Figure 2: The synergistic architecture of OOMB’s efficient KV cache management relies on four deeply interdependent core components: chunk-wise training, paged key-value cache, sparse attention, and key-value cache offload. Each optimization enables and enhances the others, collectively resolving the memory bottleneck of long-context training.
3 Preliminary: Chunk-wise Training
The core strategy of OOMB is rooted in the principles of sequential processing, which have long been used to make training Recurrent Neural Networks (RNNs) memory efficient. We first revisit this foundation and then describe how it can be adapted to the Transformer architecture, setting the stage for the challenges that OOMB is designed to solve.
Activation Recomputation in Sequential Models. In a standard RNN, the forward pass iteratively computes an output $y_{i}$ and a new hidden state $m_{i}$ at each timestep $i$:
$y_{i},m_{i},\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}a_{i}}\leftarrow\mathcal{F}_{\text{RNN}}(x_{i},m_{i-1};\Theta)$ (1)
where $\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}a_{i}}$ represents the intermediate activations cached for the backward pass. For long sequences, storing all activations $\{\mathbf{a}_{1},...,\mathbf{a}_{T}\}$ creates a memory bottleneck that scales linearly with sequence length. Activation recomputation (Sohoni et al., 2022; Bencheikh et al., 2024) circumvents this by discarding $\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}a_{i}}$ during the forward pass and regenerating it on-the-fly just before its use in the backward pass. This trades a modest amount of computation for a significant reduction in memory.
Adapting Sequential Processing to Transformers. While Transformers are inherently parallel, they can be trained using a similar chunk-wise, sequential paradigm. We partition a long input sequence into $S$ chunks, $\{X_{1},X_{2},...,X_{S}\}$, and process them serially. The forward pass for chunk $i$ attends to the KV caches from all preceding chunks, which act analogously to an RNN’s hidden state:
$Y_{i},M_{i},\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}A_{i}}=\mathcal{F}_{\text{Transformer}}(X_{i},\{M_{1},M_{2},...,M_{i-1}\};\Theta)$ (2)
Here, $M_{i}$ is the KV cache generated by chunk $i$, and $\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}A_{i}}$ represents its intermediate activations. By applying the same activation recomputation strategy, we discard $\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}A_{i}}$ after the forward step and recompute it for the backward pass:
$\text{d}\Theta,\{\text{d}M_{1},...,\text{d}M_{i-1}\}\leftarrow\text{backward}(\text{d}Y_{i},\text{d}M_{i},\mathbf{\color[rgb]{0.12109375,0.2421875,0.40234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.2421875,0.40234375}A_{i}})$ (3)
This chunk-recurrent approach ensures that the activation memory footprint remains constant, determined only by the size of a single chunk. Consequently, this paradigm fundamentally shifts the memory bottleneck: the challenge is no longer the activations, but the management of the KV cache $\{M_{1},...,M_{S}\}$ and its gradients, which still grow linearly with the sequence length.
Figure 3: Our paged KV cache management system for training. A custom kernel handles both forward and backward passes, updating gradients in-place to bypass the PyTorch autograd system for greater efficiency.
[…]
4 The OOMB Training System
OOMB builds upon the chunk-wise training paradigm by integrating three synergistic components: paged KV cache management, sparse attention, and asynchronous KV cache offloading. As illustrated in Figure 2, these components are highly interdependent, and their combined effect is critical to the system’s overall efficiency. This section provides a detailed description of each component.
4.1 Paged Memory Management for KV Cache and Gradients
Chunk-wise training requires the KV cache to grow incrementally, a process analogous to autoregressive decoding. Standard tensor concatenation for this purpose is inefficient, leading to frequent memory reallocations, copy operations, and severe memory fragmentation (Kwon et al., 2023). As shown in Figure 4, this inefficiency causes memory consumption to grow disproportionately with context length, limiting scalability.
To resolve this, OOMB incorporates a paged memory manager for both the KV cache and its gradients. Since existing paged attention solutions are designed for inference and lack backpropagation support, we developed a custom Triton kernel that handles both forward and backward passes (Figure 4).
Our implementation is designed to bypass the PyTorch autograd engine. This approach provides two key benefits. First, custom CUDA kernels accumulate gradients in-place using atomic_add, which eliminates intermediate buffers and reduces memory I/O. Second, decoupling the KV cache from the autograd graph prevents PyTorch from storing it as an activation. This design simplifies memory management and facilitates efficient offloading by maintaining a leaner computation graph.
4.2 Sparse Attention
The quadratic complexity of attention is a primary bottleneck in long-context training. OOMB addresses this by integrating page-level sparse attention, which reduces computational cost and is naturally supported by our paged KV cache architecture (Chen et al., 2024; Li et al., 2025). We implement two variants of sparse attention tailored to different model architectures.
Approximating Dense Attention for Qwen2.5. For models with dense attention like Qwen2.5 (Yang et al., 2025), we employ a Top-K page retrieval method to approximate the full attention pattern (Tang et al., 2024; Lu et al., 2025). The selected page indices from the forward pass are cached for reuse during backpropagation. The retrieval process begins by computing a single representative vector $K_{\text{avg}}\in\mathbb{R}^{n\times D}$ for each of the $n$ key pages by averaging. Given queries $Q\in\mathbb{R}^{(m\times P)\times D}$ from the current chunk, where $m$ is the number of query pages in the chunk and $P$ is the page size (tokens per page), we compute similarity scores between each query token and all page-representative vectors. A voting mechanism then aggregates these scores for each of the $m$ query pages to identify the most relevant key pages:
[…]
This retrieval process, visualized in Figure 5, incurs negligible computational overhead with page sizes ($P$) of 128 on H200 GPUs.
Native Sparse Attention for GPT-OSS. For natively sparse models like GPT-OSS (Agarwal et al., 2025), which alternate between dense and local attention layers, our implementation is straightforward. The dense layers use the same Top-K retrieval method described above, while the local attention layers retrieve only the most recent KV pages.
Figure 5: Visualization of the page-level sparse attention patterns for the Qwen2.5-7B model. Each subplot displays the retrieved key pages (x-axis) for each query page (y-axis) across the initial four layers at context lengths scaling from 32K to 256K.
4.3 Asynchronous KV Cache Offloading
The KV cache is the only component in OOMB whose memory footprint scales with sequence length. To enable training on extremely long contexts, we introduce an asynchronous offloading mechanism that transfers the KV cache between GPU and CPU memory. The offloading strategy is adapted to the attention type to maximize the overlap between data transfer and computation.
Offloading for Dense and Local Attention. For these attention patterns, the required KV pages for a given layer are known in advance. We employ a pre-fetching strategy. During the forward pass for layer i1i-1, the KV cache for layer $i$ is asynchronously fetched from the CPU. A symmetric process occurs during the backward pass. This approach effectively hides most of the data transfer latency, resulting in an end-to-end overhead lower than 5%.
Offloading for Sparse Attention. In retrieval-based sparse attention, the required KV pages are not known until the query vectors for the current layer have been computed. Consequently, during the forward pass, we initiate page retrieval and the corresponding asynchronous data transfer immediately after the query projection. The small volume of data required by sparse attention allows this transfer to be overlapped with the subsequent key and value projection computations. For the backward pass, the required page indices are already cached from the forward pass, allowing us to use the same pre-fetching strategy as in dense attention.
Both offloading schemes, illustrated in Figure 6, are implemented using pinned CPU memory and dedicated CUDA streams to ensure true asynchronicity. Data transfers are handled by the GPU’s Direct Memory Access (DMA) engine to minimize latency.
5 Experiments
Figure 6: Asynchronous KV cache offloading in OOMB. (Left) For dense and local attention, the KV cache for the next layer is pre-fetched during computation of the current layer. (Right) For sparse attention, the forward pass overlaps data transfer with key-value projections, while the backward pass reverts to the pre-fetching strategy.
5.1 Experimental Setup
Unless otherwise specified, all experiments adhere to the following configuration.
Model and Dataset. We use the Qwen2.5-7B model (Yang et al., 2025), a powerful foundation model, for all efficiency and scalability benchmarks. The training data is sourced from the arXiv dataset (Soboleva et al., 2023). To construct sequences of the required lengths for our long-context experiments, we concatenate samples from the dataset.
Baseline Methods. For efficiency evaluation, we compare our method against two baselines: (1) a standard parallel training setup using FlashAttention (Dao, 2024) and layer-wise gradient checkpointing; and (2) Ring Flash Attention (RFA) (Lin, 2025), a state-of-the-art implementation of Ring Attention (Liu et al., 2024).
Additional details are provided in Appendix B.
Table 2: Performance and Ablation of OOMB for Qwen2.5-7B on an H200 GPU. Per-iteration latency and peak memory for various OOMB configurations (ablating chunk size, token budget, gradient checkpointing, and CPU offload) and a FlashAttention (Dao, 2024) baseline. Red intensity indicates: (1) for latency, the slowdown vs. the Ckpt-only counterpart; (2) for memory, the growth rate as context length increases. Single-GPU offload results are reported under full 8-GPU system load.
Method Budget Configuration Per-iteration Latency (s) Peak Memory Usage (MB)
Ckpt. Offload 8K 16K 32K 64K 128K 256K 8K 16K 32K 64K 128K 256K
Baseline - - 2.4 5.7 20.4 76.1 OOM OOM 30298 47025 64985 1002546 OOM OOM
[…]
OOMB + Dense Attn - 2.6 6.0 19.3 68.9 260.6 1015.9 34310 36326 39804 44390 55143 76650
2.4 6.0 18.6 63.4 233.9 899.6 57369 57529 57743 58175 59065 60788
3.0 6.7 20.7 70.7 260.4 985.8 33686 33848 34076 34504 35404 37129
OOMB + Sparse Attn 4K+512 2.7 4.4 8.8 18.0 36.6 75.7 34317 36335 39826 44409 55177 76719
3.0 4.5 9.5 19.6 40.0 83.3 57348 57372 57377 57380 57410 57448
2.8 5.2 10.8 22.6 46.5 96.9 33669 33694 33999 33712 33732 33769
4K+2048 2.7 4.9 10.6 21.9 44.9 93.2 34317 36335 39827 44411 55178 76723
2.9 5.5 11.5 24.5 51.4 107.2 57348 57401 57413 57456 57479 57515
3.1 6.2 13.4 28.6 60.9 127.5 33669 33722 33735 33778 33800 33837
4K+ 8192 2.6 5.9 14.3 31.5 66.2 138.1 34317 36335 39827 44413 55183 76733
2.4 6.7 16.2 37.0 83.4 178.6 57348 57421 57487 57614 57647 57752
2.9 7.4 18.2 42.8 94.0 201.4 33669 33743 33809 33857 34083 34083
4K+ 32768 2.5 6.0 19.3 61.1 148.1 324.9 34317 36335 39826 44416 55196 76768
2.4 6.4 20.6 67.1 154.7 392.1 57348 57421 57520 57716 57716 58238
3.0 7.2 22.7 79.8 187.7 427.4 33669 33743 33842 34036 34426 34503
[…]
OOMB + Dense Attn - 2.8 5.7 19.0 68.4 259.0 1014.5 31628 40795 43483 48860 59613 81119
2.3 5.5 17.6 64.3 237.4 922.2 71025 85705 85898 86282 87051 86589
3.0 6.1 19.7 70.6 264.3 1021.4 30374 38293 38485 38870 39639 41177
OOMB + Sparse Attn 8K+512 2.5 5.1 11.4 24.1 49.6 101.9 31826 40796 43488 48872 59639 81181
2.2 5.5 12.1 25.3 52.2 107.8 70979 85602 85624 85660 85694 85726
2.5 6.3 13.4 28.7 59.9 123.7 30380 38193 38215 38253 38284 38270
8K+2048 2.5 5.4 12.4 26.3 54.4 111.6 31826 40797 43489 48873 59640 81183
2.4 5.4 12.6 27.9 58.2 120.7 70979 85610 85635 85716 85732 85757
2.6 6.6 14.7 31.9 67.3 142.1 30380 38201 38226 38262 38317 38364
8K+ 8192 2.7 5.8 16.0 36.6 78.1 162.6 31826 40797 43489 48874 59646 81192
2.4 5.7 16.3 38.7 85.2 183.4 70979 85613 85695 85813 85952 85990
3.0 6.4 17.9 41.7 94.0 204.7 30380 38204 38286 38319 38427 38608
8K+ 32768 2.8 5.7 18.9 63.1 159.8 345.6 31826 40797 43488 48877 59658 81227
2.4 5.7 18.9 64.8 165.0 355.2 70979 85613 85713 85917 86122 86122
3.0 6.4 20.8 70.8 177.7 408.5 30380 38204 38304 38508 38731 38856
[…]
OOMB + Dense Attn - 2.8 6.1 19.6 70.9 262.5 1024.0 31627 44589 46605 52654 62735 84913
2.3 5.7 18.9 68.2 258.7 1017.2 71306 113827 115169 114398 115190 117984
2.8 6.1 19.6 70.9 262.5 1024.0 31575 42705 42879 44081 45858 45858
OOMB + Sparse Attn 12K+512 2.7 4.7 12.5 28.5 60.5 125.3 31626 44589 46609 52064 62760 84973
2.4 4.7 12.3 28.6 60.3 126.3 70979 113753 113765 113810 113881 113890
2.9 5.4 14.0 32.8 72.8 146.4 30379 42639 42681 42698 42737 42774
12K+2048 2.8 4.9 13.5 31.3 66.7 138.1 31626 44589 46609 52065 62761 84975
2.4 5.1 13.9 32.8 69.8 146.7 70979 113753 113817 113880 113923 113963
3.0 5.7 15.2 36.4 78.9 163.5 30379 42609 42703 42766 42809 42845
12K+ 8192 2.8 5.7 17.6 43.4 90.6 188.7 31626 44593 46609 52067 62765 84985
2.4 5.8 18.4 43.6 90.5 197.9 70979 113753 113830 113997 114017 114138
3.0 6.6 21.5 48.2 105.1 224.1 30379 42639 42716 42860 42968 43018
12K+ 32768 2.8 5.7 18.9 66.4 170.0 373.9 31626 44589 46609 52070 62778 85019
2.4 5.7 18.5 67.2 169.0 385.3 70979 113753 113809 114028 114284 114558
3.1 6.5 20.6 74.0 185.6 421.2 30379 42639 42715 42944 43204 43443
5.2 Memory, Time, and Scalability Analysis
We evaluate OOMB’s efficiency against two paradigms: standard parallel training and context parallelism.
Comparison with Parallel Training. Table 2 presents a comprehensive efficiency comparison between OOMB and a parallel training baseline using FlashAttention. Our analysis highlights several key findings:
  • At a 32K context length, OOMB with dense attention already outperforms the baseline, and this performance advantage widens as the context length increases.
  • The asynchronous offloading mechanism imposes negligible latency overhead on dense attention, demonstrating its high efficiency.
[…]
  • For sparse attention with offloading, a larger chunk size mitigates latency; a 12K chunk size, for example, adds only a 12% overhead.
  • Our offloading system dramatically improves memory scalability for both dense and sparse attention, keeping peak memory usage nearly constant irrespective of context length.
  • Even without CPU offloading, OOMB reduces the memory growth rate by more than fivefold compared to the parallel training baseline.
Table 3: Per-device training throughput (tokens/sec) comparison against context-parallel methods. All OOMB configurations use a 4K chunk size with checkpointing and CPU offload; the sparse attention version adds an 8192 retrieval budget. TP: Tensor Parallel.
Context Ring Attn (report) RFA OOMB (Dense Attn) OOMB (Sparse Attn)
[…]
64K - - 872.51 936.22 933.84 1560.38 1542.85
128K - - 463.37 504.12 500.26 1394.16 1371.41
256K 50 49 218.41 266.13 261.47 1301.60 1265.59
Comparison with Context Parallelism Methods. We also benchmarked OOMB against Ring Flash Attention (RFA) (Lin, 2025), a state-of-the-art implementation of Context Parallelism. This paradigm manages memory by splitting sequences across GPUs but introduces significant inter-device communication overhead. As shown in Table 3, OOMB achieves higher per-device training throughput, outperforming the heavily optimized RFA.
Figure 7: (Top-left) Training loss curves for Qwen2.5-7B (Yang et al., 2025) with 256K and 1M context lengths, employing sparse attention with retrieval budgets of 4K+8192 and 4K+32768. (Top-right) Distribution of the L2 norm of gradient approximation error for sparse attention across various context lengths (16K to 1M) and retrieval budgets. (Bottom) Layer-wise L2 norm of the gradient estimation error for different projection matrices, analyzed with a 64K context length and a 4K+8192 retrieval budget. Our work offers a preliminary validation of this sparse training; for more optimal sparse attention strategies, we refer readers to the established literature.
5.3 Accuracy Validation of Sparse Attention
We empirically validated the accuracy of our sparse attention mechanism and its impact on model training. First, we analyzed the gradient approximation error for Qwen2.5-7B (Yang et al., 2025) across 16 configurations, composed of 4 context lengths and 4 retrieval budgets. As shown in Figure 7 (Top-right), within the model’s 128K native context limit, a larger retrieval budget results in a smaller approximation error. Beyond this limit, the error differences between budgets become less significant. The loss values for these settings, presented in Table 4(a), further indicate that sparse attention does not substantially degrade model performance.
[…]
Collectively, these findings confirm that sparse attention is a viable and effective method for long-context training, as it preserves model performance while achieving a constant $\mathcal{O}$(1) memory footprint.
Table 4: (a) Training loss for the 16 configurations (4 context lengths $\times$ 4 sparse budgets) in the top-right subplot of Figure 7. (b) Resource usage during training for the 4 configurations (2 context lengths $\times$ 2 sparse budgets) in the top-left subplot of Figure 7. These experiments were conducted on four H200 GPUs using tensor parallelism, without gradient checkpointing or CPU offload.
(a) Loss values under different settings
Context Dense Attention Sparse Attention
4K+512 4K+2048 4K+8192 4K+32768
16K 1.32031 1.50000 1.36718 1.32031 1.32031
64K 1.22387 1.47656 1.32812 1.21875 1.21875
256K 2.21875 2.34375 2.28125 2.21875 2.21875
1M 4.31250 4.31250 4.50000 4.71875 4.78125
[…]
Metric 4K+8192 4K+32768
256K 1M 256K 1M
Latency/Iter (s) 43.454 176.305 87.340 422.383
Peak Memory (MB) 35637 67944 35671 68145
6 Conclusion and Limitations
We introduced OOMB, a highly memory-efficient training system that overcomes the primary memory barrier in long-context LLM training. By integrating a chunk-recurrent framework with activation recomputation, a paged KV cache manager, asynchronous CPU offloading, and page-level sparse attention, OOMB maintains a nearly constant memory footprint regardless of sequence length. Our results demonstrate that OOMB enables training a Qwen2.5-7B model with a 4-million-token context on a single H200 GPU, a task that would otherwise require a large-scale cluster. This work represents a significant advance in making long-context model training more accessible and resource-efficient.
Despite its efficiency, OOMB has several limitations. The chunk-wise serial processing introduces a latency overhead compared to fully parallel training paradigms, which may be a consideration for throughput-sensitive training scenarios. Additionally, our use of sparse attention is an approximation of the full attention mechanism, and its impact on model performance for tasks requiring dense global context requires further investigation. The effectiveness of the asynchronous offloading is also dependent on the available CPU-GPU interconnect bandwidth.
Ethics Statement
The primary contribution of this work is a system designed to improve the computational efficiency of training large language models on long contexts. By significantly reducing the hardware requirements for such tasks, our research aims to democratize access to long-context model development for researchers and institutions with limited computational resources. A direct positive implication of this efficiency is the potential reduction in the overall energy consumption and carbon footprint associated with training these models, promoting more sustainable AI research practices.
We acknowledge the potential for dual use, as is common with foundational AI research. The tools we develop could be applied to train models for malicious purposes, such as generating misinformation. While this work focuses on the system’s technical aspects, we advocate for the responsible development and deployment of models trained using our framework. Our experiments were conducted on the publicly available arXiv dataset, which consists of scientific literature and does not contain personally identifiable or sensitive information.
Reproducibility Statement
To facilitate the reproducibility of our findings, we have made the complete source code for the OOMB training system available in the anonymous repository cited in the abstract. A detailed description of our experimental setup, including model specifications, datasets, hyperparameters, and evaluation protocols, is provided in Section 5.1 and Appendix B. The core methodology and its components are thoroughly explained in Section 4. We believe these resources offer a sufficient basis for researchers to verify our results and extend our work.
Appendix A Use of Large Language Models (LLMs)
During the writing of this paper, we utilized LLM solely for language editing to improve clarity and readability. We critically reviewed and revised all AI-generated suggestions to ensure the final text accurately reflects our original intent. All intellectual contributions, including the research design, methodology, analysis, and conclusions, are our exclusive work, and we take full responsibility for the academic integrity of this publication.
Appendix B Experimental Details
All experiments were conducted using full-parameter fine-tuning in bfloat16 precision. We employed the Adam optimizer with a learning rate of $5\times 10^{-5}$ and betas of $(0.9,0.98)$. To isolate the impact of sequence length on system performance, the per-GPU batch size was set to one, and all other model and training hyperparameters were maintained at their default settings. For OOMB-specific configurations, the KV cache page size was set to 128 for H200 GPUs.
System performance was evaluated using two primary metrics: peak GPU memory and per-iteration latency. Peak memory per device was recorded using PyTorch’s max_memory_allocated() function, while latency was measured precisely with CUDA events. To ensure reliability and mitigate measurement noise, all reported results represent the minimum value from three independent runs.
Appendix C Additional Experiments and Ablation Studies
This section provides supplementary experiments that further validate the performance, scalability, and correctness of the OOMB framework.
Performance and Ablation Studies
Scalability to 4M+ Tokens. A key objective of OOMB is to enable training on sequence lengths that are infeasible with conventional methods on limited hardware. We benchmarked the per-iteration training time and memory usage of OOMB on contexts up to 8 million tokens. Figure 8 shows that OOMB maintains exceptional memory efficiency even at this scale. When using sparse attention, the training time scales in a near-linear fashion, making ultra-long context training practically achievable.
[…]
Ablation on Chunk Size. The chunk size is a key hyperparameter in OOMB that balances computational parallelism and activation memory. We evaluated the effect of varying the chunk size from 512 to 4096. As shown in Figure 9, larger chunk sizes improve computational efficiency by enabling better hardware utilization. However, these performance gains diminish beyond a size of 4096. Memory consumption remains stable across all tested sizes. This analysis supports 4096 as a well-balanced default value, as it provides strong computational performance without excessive memory usage for activations.
Figure 9: Ablation on Chunk Size. Increasing the chunk size improves computational throughput with diminishing returns, while peak memory usage remains stable. A chunk size of 4096 offers a strong balance between efficiency and memory. This experiment was conducted on a single A100 80G GPU.